Patrick spent 4/9 of an hour (h) in the batting cages on Friday and 7/9 of an hour on Saturday. He wants to spend 2 h total on the weekend. How much time should he spend on Sunday to accomplish this goal?

Answer by mananth(16946)

4/9+7/9=11/9 he has spent
Balance 2-11/9 = 18-11 / 9
=7/9 hours
He has 7/9 hours for sunday